graph the equation by plotting three points : 2y = -x + 2

Answer by mananth(16949)

1 x + 2 y = 2
y= - 1/ 2 x + 1

Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 4 2
y= 1 -1 0
(x,y)(0,1),(4,-1),(2,0)